# Self-Contrast: Better Reflection Through Inconsistent Solving Perspectives

### v1

@@ Abstract
[…]
@@ 1 Introduction
As a remedy, we propose a contrastive strategy as an alternative to the direct evaluation: we examine the differences among multiple responses and draw inspiration to derive feedbacks from their disparities for reflection. The insight is that while generating accurate feedback directly may be challenging, identifying contrasts between various responses is often more feasible. More importantly, these discrepancies often indicate some potential errors, easily overlooked details or pitfalls. As shown in Figure 1, by contrasting two solutions, LLM finds they have different solving objectives, and suggests re-examining the intent of the original request in the checklist.
@@ 2.2 Feedback Analysis
Fail to Correct the Wrong Initial Response. As shown in Table 2, we observe the number of Toxic Reflection (normal-$\Rightarrow$: 52) and Valid Reflection (normal-$\Rightarrow$: 48) are nearly similar. This explains why there is no discernible difference in performance pre- and post-reflection. Besides, considering the scenario when the initial response is erroneous, we observe the number of Invalid Reflection (normal-$\Rightarrow$: 269) is significantly larger than Valid Reflection (normal-$\Rightarrow$: 48), which indicates LLM fails to correct errors in the initial responses for most cases.
@@ 2.3 From Self-Evaluate to Self-Contrast
Table 3: We report the accuracy change ($\Delta$) between post- and pre-reflection for 3 settings and t-test value for $\Delta$¿0. Self-evaluate: Directly evaluate the initial response. Self-contrast: Contrast the difference between two responses and generate a checklist for reflection.
@@ 3 Self-Contrast
As shown in Figure 2, LLM designs five different prompts and their translation results based on the user’s request (”这个计划被枪毙”) . From a literal perspective, the phrase ”被枪毙” is translated as ”shot to death”. This rigid translation fails to grasp the metaphor embedded in the military term. Conversely, from a liberal perspective, it is translated as ”This plan was axed”. After contrasting two different translations, LLMs believe they should scrutinize the source sentence for metaphors and ensure the translation aligns with the conventions of English expression.
@@ 3.1 Create Diverse Perspectives
Self-Curated Prompts First, it is imperative to define the concept of ”solving perspective”. It refers to deliberate prompting with a unique role, personality, thought style, etc., which prompts LLMs to solve user requests from a specific perspective. Diverse solving perspectives can endow LLMs with a broader range of thoughts for problem-solving, e.g., different angles and methodologies, thereby mitigating biases introduced by singular prompts.
@@ 3.2 Contrast Inter-Perspective Discrepancies
Contrasting After selecting $k$ responses from all candidates, we feed these responses concurrently into LLM and then instruct LLM to autonomously contrast the differences for each pair of responses in a single pass. When contrasting, LLMs need to explicitly answer these questions: Whether the two responses are different, Where the differences lie, and Which factors contributed to these inconsistent results. These questions guide the LLM to methodically explore the underlying reasons behind discrepancies, identifying potential errors and often overlooked details. As shown in Figure 2, for translation tasks, the LLM compares results 1, 2, and 5, and identifies that their primary differences lie in the use of different verbs to express ”被枪毙”. The detailed prompts are shown in Section D.2.
@@ 4 Experiments
Self-Contrast 84.4 $\uparrow$7.8 95.4 $\uparrow$1.5 20.5 $\uparrow$0.7 42.3 normal-$\uparrow$9.2 64.2 $\uparrow$11.6 89.0 $\uparrow$9.2 94.0 $\uparrow$1 44.5 $\uparrow$7 54.6 $\uparrow$14.4 75.3 $\uparrow$9.3 7.8
Table 4: The performance on mathematical reasoning. For a fair comparison, Self-Consistency (SC-Vote, -Select, -Reflect) samples eight responses and then performs voting, selecting, or reflection. For the Multi-Agent, we configure three agents to engage in a three-round debate. $\uparrow$ and $\downarrow$ means accuracy changes over the CoT prompt.
@@ 4.1 Settings
Benchmarks We evaluate our method within two challenging testbeds: mathematical reasoning and creative translation using GSM8K, SVAMP, and CommonMT benchmarks. Please refer to Section B.1 for a detailed description.
@@
LLM Models and Prompts We conduct experiments using the GPT-3.5-Tubor-0613 and GPT-4-0613 APIs, alongside the Llama2-Chat model with three parameter scales (7B, 13B, and 70B). To make a fair comparison, we uniformly set the temperature to 0.2 for all experiments. For standard prompts and self-reflection baseline, we evaluate them 10 times using different prompts and average their results under zero-shot scenes. Details of the prompts and other parameters can be found in Appendices C, D and B.2.
@@ 4.3 Main Results
Better generality across different LLMs and tasks. From widely used commercial LLMs (e.g., GPT4) to open-source models (Llama-2), and from reasoning to generative tasks, our strategy exhibits robust generalizability. Concretely, from the perspective of LLM, Self-Contrast achieves the best results on most models except Llama-2-7B. For instance, for GPT-3.5, there is an enhancement of 7.8% on GSM8K and 9.2% on SVAMP, while for Llama-2-70B, the improvements are 11.6% and 9.3% respectively. As for Llama-2-7B, our performance is only slightly lower than Self-consistency and Multi-Agent. We hypothesize this might be due to the weaker instruction-following capabilities of the Llama2-7B, making it challenging to accurately contrast two inconsistent solutions. Besides task-wise, Self-Contrast applies to various task types, demonstrating high versatility. In contrast, Self-Consistency can not handle non-numerical tasks directly, e.g., translation, due to its voting mechanism (Table 5). Its variant strategies, SC-Select and SC-Reflect, lag significantly behind ours.
Fewer manual efforts and more reasonable call overheads. Compared to the multi-agent debate, Self-Contrast gains more significant improvements with less call overhead (¿10% reduction). From a unified perspective, Self-Contrast can be viewed as a multi-agent contrastive mechanism. Instead of a free-form debate among multiple agents, our strategy fosters a more explicit and purposeful debate by contrasting the differences between agents and summarizing the reasons for their disagreements. Moreover, Self-Contrast is flexible, dynamically designing multiple perspectives in response to user requests, without the need for manually pre-configuring agent roles and quantities.
@@ 5 The Effect of the Different Components
Self-curated Prompt Vs. Sampling Multiple Responses. Instead of self-curated prompt process, we directly sample multiple responses from LLMs for subsequent contrast and reflection. As shown in Figure A2, we observe that as the number of sampled responses increases, the final performance also improves, yet it still remains lower than Self-Contrast with the self-curated prompts process, where full strategy achieves 84.4% compared to the maximum of 81.8% when sampling 5 responses. We find that the top-n responses are sometimes strikingly similar, diminishing the effectiveness of the contrastive strategy.
Reflection Without Checklist. Next, we eliminate the checklist generation process, i.e., directly instruct the LLM to reflect on the differences among perspectives. As shown in Table A1, we find it brings a significant impact on mathematical reasoning (-3.5%), but a slight impact on translation (-0.1%). This is because translation tasks tend to focus more on local features. Even without a checklist, the LLM can also effectively reflect based on the comparisons of lexical, syntactic.
@@ 6 Analysis
LLMs Strategy Invalid Cases $\downarrow$ Toxic Cases $\downarrow$
@@
Table 6: Self-Contrast is evaluated on two cases. Invalid: Wrong $\rightarrow$ Wrong and Toxic: Right $\rightarrow$ Wrong.
6.1 Self-Contrast Reduces Invalid and Toxic Reflection
@@
6.2 Contrasting Two Incorrect Solutions is also Instructive
Self-Contrast inspires reflection by contrasting the differences. An intuitive explanation is that the errors in different responses are dissimilar or randomized, so that they can be used to compare with each other and eliminate uncertainties or biases. To verify this, we sample 200 questions from GSM8K, each manually annotated with a correct solution, two incorrect solutions containing similar errors (e.g., Error1), and an incorrect solution with a different error (Error2). We design four experiments: 1. Self-evaluate one incorrect solution followed by reflection. 2. Self-Contrast a correct and an incorrect solution. 3. Self-Contrast two similar incorrect solutions. 4. Self-Contrast two dissimilar incorrect solutions. As shown in Table 7, we observe that contrasting a correct and an erroneous solution, or contrasting two incorrect solutions with different errors both yield significant enhancements of 13.5% and 5.4%, respectively. In contrast, comparing two solutions that contain similar errors does not result in perceptible changes.
@@ B.1 Benchmarks
Creative Translation In addition to mathematical reasoning tasks, we introduce a generation task: creative translation. We utilize the CommonMT He et al. (2020), which includes a vast body of Chinese-to-English pair examples. Unlike conventional translations, most samples contain non-standard expressions such as idioms and metaphors, necessitating an understanding of local cultural and linguistic habits for accurate translation. Following the Multi-agent debate (Liang et al., 2023), we adopt the samples with ”hard” categories from CommonMT as testing benchmarks.
@@ D.1 Prompt for Self-Curated Process
The input Chinese sentence is: 他想拉同村的干部一起下水去贩毒. Please generate the most suitable prompts.
@@ E.2 Prompting for Better Problem-Solving
Drawing on cognitive science, human reasoning involves two different reasoning patterns: breadth reasoning, i.e., exploring various reasoning perspectives, and depth reasoning, which involves continually refining ideas and minimizing errors. Based on this concept, we can view previous prompting strategies as either breadth or depth reasoning. Self-consistency and some contemporaneous works (Wang et al., 2023d; Huang et al., 2022; Yoran et al., 2023; Jain et al., 2023; Chen et al., 2023c) mimic breadth reasoning by sampling diverse reasoning processes and voting the final answer, while self-reflection, abstraction reasoning strategies (Shinn et al., 2023; Madaan et al., 2023; Paul et al., 2023; Zheng et al., 2023a; Wang et al., 2023a; Yoran et al., 2023; Zheng et al., 2023b; Xu et al., 2023b) represent depth reasoning, refining reasoning through iterative prompting strategy. Except for these, Self-Verification (Weng et al., 2022) designs a reverse generation from the answer to given conditions, which is widely used in machine translation (Edunov et al., 2018), reinforcement learning (Zhang et al., 2021, 2022c). Cohen et al. (2023); Mündler et al. (2023) propose a method for detecting self-contradictions or factual errors in responses to enhance quality. However, our Self-Contrast combines both breadth and depth reasoning. It creates multiple perspectives to enhance the breadth of reasoning and also reflects on the differences for better depth reasoning, offering more reliable problem-solving.
@@ E.4 Learning Mathematical Reasoning
In recent years, mathematical reasoning has become a significant benchmark (Cobbe et al., 2021; Hendrycks et al., 2021) to evaluate the capabilities of artificial intelligence models. Within the paradigm of supervised learning, a vast amount of research (Xie and Sun, 2019; Patel et al., 2021; Jie et al., 2022; Zhang et al., 2022b, 2023b) has been dedicated to translating human language into mathematical equations. In the era of LLMs, the advent of Chain-of-Thought and other prompting strategies have notably augmented the reasoning capabilities (Zhu et al., 2023; Yuan et al., 2023b; Frieder et al., 2023; Zhou et al., 2022).
Prompting Method PAL and Program-of-Thoughts Gao et al. (2022); Chen et al. (2022) separate the computation and reasoning process using code as the intermediate process. Mathprompter, Auto-Model (Imani et al., 2023; Zhao et al., 2023) encourage LLMs to generate diverse reasoning paths in different forms simultaneously, including text (CoT), code (PAL), and symbols (Equation) for a higher confidence answer. Automatic-CoT, Complexity-CoT, Synthetic Prompt and Boosted Prompt (Zhang et al., 2022d; Fu et al., 2022; Shao et al., 2023; Pitis et al., 2023) enhance reasoning performance by optimizing the selection of demonstrations within the prompt. Tree-of-thought and Self-Evaluation (Yao et al., 2023; Xie et al., 2023) extend the CoT into a search tree, obtaining more accurate answers through self-evaluation.

### v2

@@ Abstract
Wenqi Zhang1, Yongliang Shen1, Linjuan Wu1 Qiuying Peng2, Jun Wang2, Yueting Zhuang1, Weiming Lu1 1College of Computer Science and Technology, Zhejiang University 2OPPO Research Institute, China {zhangwenqi, luwm}@zju.edu.cn
@@ 1 Introduction
As a remedy, we propose a contrastive strategy as an alternative to the direct evaluation: we examine the differences among multiple responses and draw inspiration to derive feedbacks from their disparities for reflection. The insight is that while generating accurate feedback directly may be challenging, identifying contrasts between various responses is often more feasible. More importantly, these discrepancies often indicate some potential errors, easily overlooked details or pitfalls. As shown in Figure 1, by contrasting two solutions, LLM finds they have different solving objectives, and suggests re-examining the intent of the original request in the checklist. This contrasting paradigm can also be seen in some contemporaneous work (Wan et al., 2023; Yuan et al., 2024).
@@ 2.2 Feedback Analysis
Fail to Correct the Wrong Initial Response. As shown in Table 2, we observe the number of Toxic Reflection ($\Rightarrow$: 52) and Valid Reflection ($\Rightarrow$: 48) are nearly similar. This explains why there is no discernible difference in performance pre- and post-reflection. Besides, considering the scenario when the initial response is erroneous, we observe the number of Invalid Reflection ($\Rightarrow$: 269) is significantly larger than Valid Reflection ($\Rightarrow$: 48), which indicates LLM fails to correct errors in the initial responses for most cases.
@@ 2.3 From Self-Evaluate to Self-Contrast
Table 3: We report the accuracy change ($\Delta$) between post- and pre-reflection for 3 settings and t-test value for $\Delta$>0. Self-evaluate: Directly evaluate the initial response. Self-contrast: Contrast the difference between two responses and generate a checklist for reflection.
@@ 3 Self-Contrast
As shown in Figure 2, LLM designs five different prompts and their translation results based on the user’s request {CJK*}UTF8gkai("这个计划被枪毙") . From a literal perspective, the phrase {CJK*}UTF8gkai"被枪毙" is translated as "shot to death". This rigid translation fails to grasp the metaphor embedded in the military term. Conversely, from a liberal perspective, it is translated as "This plan was axed". After contrasting two different translations, LLMs believe they should scrutinize the source sentence for metaphors and ensure the translation aligns with the conventions of English expression.
@@ 3.1 Create Diverse Perspectives
Self-Curated Prompts First, it is imperative to define the concept of "solving perspective". It refers to deliberate prompting with a unique role, personality, thought style, etc., which prompts LLMs to solve user requests from a specific perspective. Diverse solving perspectives can endow LLMs with a broader range of thoughts for problem-solving, e.g., different angles and methodologies, thereby mitigating biases introduced by singular prompts.
@@ 3.2 Contrast Inter-Perspective Discrepancies
Contrasting After selecting $k$ responses from all candidates, we feed these responses concurrently into LLM and then instruct LLM to autonomously contrast the differences for each pair of responses in a single pass. When contrasting, LLMs need to explicitly answer these questions: Whether the two responses are different, Where the differences lie, and Which factors contributed to these inconsistent results. These questions guide the LLM to methodically explore the underlying reasons behind discrepancies, identifying potential errors and often overlooked details. As shown in Figure 2, for translation tasks, the LLM compares results 1, 2, and 5, and identifies that their primary differences lie in the use of different verbs to express {CJK*}UTF8gkai"被枪毙". The detailed prompts are shown in Section D.2.
@@ 4 Experiments
Self-Contrast 84.4 $\uparrow$7.8 95.4 $\uparrow$1.5 20.5 $\uparrow$0.7 42.3 $\uparrow$9.2 64.2 $\uparrow$11.6 89.0 $\uparrow$9.2 94.0 $\uparrow$1 44.5 $\uparrow$7 54.6 $\uparrow$14.4 75.3 $\uparrow$9.3 7.8
Table 4: The performance on mathematical reasoning. Self-Consistency (SC-Vote, -Select, -Reflect) samples eight responses and then performs voting, selecting, or reflection. For the Multi-Agent, we configure three agents to engage in a three-round debate. $\uparrow$ and $\downarrow$ means accuracy changes over the CoT prompt. L- denotes Llama2-chat.
@@ 4.1 Settings
Benchmarks We evaluate our method within two testbeds: mathematical reasoning and translation using GSM8K, SVAMP, and CommonMT benchmarks. Please see Section B.1 for details.
@@
LLM Models and Prompts We conduct experiments using the GPT-3.5-Tubor-0613 and GPT-4-0613, alongside the Llama2-Chat model with three parameter scales (7B, 13B, and 70B). To make a fair comparison, we uniformly set the temperature to 0.2 for all experiments. For standard prompts and self-reflection baseline, we evaluate them 10 times using different prompts and average their results under zero-shot scenes. Prompts and other details can be found in Appendices C, D and B.2.
@@ 4.3 Main Results
Better generality across different LLMs and tasks. From commercial LLMs (e.g., GPT4) to open-source models (Llama-2), and from reasoning to generative tasks, our strategy exhibits robust generalizability. Concretely, from the perspective of LLM, Self-Contrast achieves the best results on most models except Llama-2-7B. For instance, for GPT-3.5, the improvements are 7.8% on GSM8K and 9.2% on SVAMP, while for Llama-2-70B, the improvements are 11.6% and 9.3% respectively. As for Llama-2-7B, our performance is slightly lower than Self-consistency and Multi-Agent. This might be due to the weaker instruction-following capabilities of the Llama2-7B, making it challenging to contrast two inconsistent solutions. Besides task-wise, Self-Contrast applies to various task types, demonstrating high versatility. In contrast, Self-Consistency can not handle non-numerical tasks directly, e.g., translation, due to its voting mechanism (Table 5). Its variant strategies, SC-Select and SC-Reflect, lag significantly behind ours.
Fewer manual efforts and more reasonable call overheads. Compared to the multi-agent debate, Self-Contrast gains more significant improvements with less call overhead (>10% reduction). From a unified perspective, it can be viewed as a multi-agent contrastive mechanism. Instead of a free-form debate among multiple agents, our strategy fosters a more explicit and purposeful debate by contrasting the differences between agents and summarizing the reasons for their disagreements. Moreover, Self-Contrast is flexible, dynamically designing multiple perspectives tailored to user requests, without the need for manually pre-configuring agent roles and quantities.
@@ 5 The Effect of the Different Components
Self-curated Prompt Vs. Sampling Multiple Responses. Instead of self-curated prompt process, we directly sample multiple responses from LLMs for subsequent contrast and reflection. Figure A2 shows that the final accuracy improves as the number of sampled responses increases, yet it is still lower than Self-Contrast with self-curated prompts process, where full strategy achieves 84.4% compared to the maximum of 81.8% when sampling 5 responses. We find that the top-n responses are sometimes strikingly similar, diminishing the effectiveness of the contrastive strategy.
Reflection Without Checklist. We eliminate the checklist generation process, i.e., directly instruct the LLM to reflect on the differences among perspectives. In Table A1, it brings a significant impact on mathematical reasoning (-3.5%), but a slight impact on translation (-0.1%), since translation tasks tend to focus more on local features. Even without a checklist, the LLM also can reflect based on the comparisons of lexical, syntactic.
@@ 6 Analysis
LLMs Strategy Invalid $\Rightarrow$ Cases $\downarrow$ Toxic $\Rightarrow$ Cases $\downarrow$
@@
Table 6: Self-Contrast is evaluated on two cases.
6.1 Reducing Invalid and Toxic Reflections
@@
6.2 Contrasting Incorrect Solutions is also Instructive
Self-Contrast inspires reflection by contrasting the differences. An intuitive explanation is that the errors in different responses are dissimilar or randomized, so they can be used to compare with each other and eliminate uncertainties or biases. To verify this, we sample 200 questions from GSM8K, each manually annotated with a correct solution, two incorrect solutions with similar errors (e.g., Error1), and an incorrect solution with a different error (Error2). We design four experiments: 1. Self-evaluate one incorrect solution followed by reflection. 2. Self-Contrast a correct and an incorrect solution. 3. Self-Contrast two similar incorrect solutions. 4. Self-Contrast two dissimilar incorrect solutions. Table 7 shows that contrasting a correct and an erroneous solution, or contrasting two incorrect solutions with different errors both yield significant enhancements of 13.5% and 5.4%. In contrast, comparing two solutions with similar errors does not result in perceptible changes.
@@ B.1 Benchmarks
Creative Translation In addition to mathematical reasoning tasks, we introduce a generation task: creative translation. We utilize the CommonMT He et al. (2020), which includes a vast body of Chinese-to-English pair examples. Unlike conventional translations, most samples contain non-standard expressions such as idioms and metaphors, necessitating an understanding of local cultural and linguistic habits for accurate translation. Following the Multi-agent debate (Liang et al., 2023), we adopt the samples with "hard" categories from CommonMT as testing benchmarks.
@@ D.1 Prompt for Self-Curated Process
The input Chinese sentence is: {CJK*}UTF8gkai他想拉同村的干部一起下水去贩毒. Please generate the most suitable prompts.
@@ E.2 Prompting for Better Problem-Solving
Drawing on cognitive science, human reasoning involves two different reasoning patterns: breadth reasoning, i.e., exploring various reasoning perspectives, and depth reasoning, which involves continually refining ideas and minimizing errors. Based on this concept, we can view previous prompting strategies as either breadth or depth reasoning. Self-consistency and some contemporaneous works (Wang et al., 2023d; Huang et al., 2022; Yoran et al., 2023; Jain et al., 2023; Chen et al., 2023c) mimic breadth reasoning by sampling diverse reasoning processes and voting the final answer, while self-reflection, abstraction reasoning strategies (Shinn et al., 2023; Madaan et al., 2023; Paul et al., 2023; Zheng et al., 2023a; Wang et al., 2023a; Yoran et al., 2023; Zheng et al., 2023b; Xu et al., 2023b; Shridhar et al., 2023) represent depth reasoning, refining reasoning through iterative prompting strategy. Except for these, Self-Verification (Weng et al., 2022) designs a reverse generation from the answer to given conditions, which is widely used in machine translation (Edunov et al., 2018). Cohen et al. (2023); Mündler et al. (2023) propose a method for detecting self-contradictions or factual errors in responses to enhance quality. However, our Self-Contrast combines both breadth and depth reasoning. It creates multiple perspectives to enhance the breadth of reasoning and also reflects on the differences for better depth reasoning, offering more reliable problem-solving.
@@ E.4 Learning Mathematical Reasoning
In recent years, mathematical reasoning has become a significant benchmark (Cobbe et al., 2021; Hendrycks et al., 2021) to evaluate the capabilities of artificial intelligence models. Within the paradigm of supervised learning, a vast amount of research (Xie and Sun, 2019; Patel et al., 2021; Jie et al., 2022; Zhang et al., 2023b) has been dedicated to translating human language into mathematical equations. In the era of LLMs, the advent of Chain-of-Thought and other prompting strategies have notably augmented the reasoning capabilities (Zhu et al., 2023; Yuan et al., 2023b; Frieder et al., 2023; Zhou et al., 2022).
Prompting Method PAL and Program-of-Thoughts Gao et al. (2022); Chen et al. (2022) separate the computation and reasoning process using code as the intermediate process. Mathprompter, Auto-Model (Imani et al., 2023; Zhao et al., 2023) encourage LLMs to generate diverse reasoning paths in different forms simultaneously, including text (CoT), code (PAL), and symbols (Equation) for a higher confidence answer. Automatic-CoT, Complexity-CoT, Synthetic Prompt and Boosted Prompt (Zhang et al., 2022b; Fu et al., 2022; Shao et al., 2023; Pitis et al., 2023) enhance reasoning performance by optimizing the selection of demonstrations within the prompt. Tree-of-thought and Self-Evaluation (Yao et al., 2023; Xie et al., 2023) extend the CoT into a search tree, obtaining more accurate answers through self-evaluation.